# Systolic Arrays and Structured Pruning Co-design for Efficient Transformers in Edge Systems

### v1

@@ Abstract
Efficient deployment of resource-intensive transformers on edge devices necessitates cross-stack optimization. We thus study the interrelation between structured pruning and systolic acceleration, matching the size of pruned blocks with the systolic array dimensions. In this setting, computations of pruned weight blocks can be skipped, reducing run-time and energy consumption, but potentially impacting quality of service (QoS). To evaluate the trade-offs between systolic array size and sparsity opportunities, we present a novel co-design framework that integrates algorithmic optimization, system simulation, and hardware design. Targeting speech recognition using transformers as case study, we analyze how configuration choices across the stack affect performance metrics. Results demonstrate that structured pruning on systems featuring systolic array acceleration can effectively increase performance, while maintaining high QoS levels. Up to 26% system-wide speedups due to structured pruning were measured, with only 1.4% word error rate degradation on the standard Librispeech dataset.
Index Terms:
@@
Systolic array, structured pruning, hardware-software co-design, edge AI.
I Introduction
@@
Transformers have fostered a revolution in machine learning, with applications ranging from classification [1] to generative models for text and images [2], to speech recognition [3]. However, their complex structure based on multiple attention and feed-forward layers [4] results in unprecedented computational requirements, posing significant challenges for their deployment. These are particularly acute in edge scenarios, where systems have to operate within constrained energy and performance envelopes.
@@
In this context, a plethora of optimization strategies have been proposed. On the software side [5], commonly used approaches involve reducing the precision of data representations (quantization) and removing parts that contribute the least to inference outcomes (pruning). As for hardware, efforts have mainly focused on the acceleration of the main computational kernel in transformers, i.e. General Matrix Multiplications (GEMMs). Although diverse solutions ranging from analog crossbars [6][7] to near-DRAM computing [8][9] work toward this goal, a particularly promising alternative is represented by systolic arrays [10]. These two-dimensional meshes of processing elements can indeed parallelize the computation of a GEMM (or, more precisely, the computation of a GEMM tile), while presenting a high parallelism degree, low resource requirements and only mandating a simple, low-overhead control logic.
@@
Recent works [11, 12, 13, 14, 15] have attempted to co-optimize software algorithms and hardware accelerators dedicated to transformer inference [16]. Such a stance is particularly appealing at the crossroads of model pruning and systolic array acceleration. On the software side, pruning can be performed by eliding weights in regular block patterns (in a “structured” way) rather than as individual elements [17]. While this approach introduces a constraint to pruning, and can hence result in lower overall sparsity rates, it substantially amplifies hardware-side optimization opportunities when matching the sizes of the pruned tile and the accelerator mesh. The exploration of this strategy, which we term Systolic Array Structured Pruning (SASP), is the focus of this work.
@@
To explore these interrelations, we employ a holistic approach integrating methods for a) the structured pruning of transformer algorithms, b) the system-level level modeling of accelerated systems executing them, and c) the hardware synthesis of accelerators. Our environment for SASP exploration builds on frameworks for the training of transformers (ESPnet [3]) and for system simulation (gem5 [18]). By employing a novel systolic array architectural template, it supports both floating point and weight-quantized data representations, as supported by ESPnet.
As a test case, we employ our exploration approach to analyze a speech recognition application, based on an 24-block, 75M-parameter transformer processing the LibriSpeech dataset [19]. We observed that SASP can achieve, for a systolic array size of $32\times 32$, up to 44% speedup and 42% energy savings over a non-pruned, non-quantized system when employing a 20% pruning rate, resulting in a marginal Word Error Rate (WER) degradation of 1.4%.
@@
  • We introduce a methodology for the systematic exploration of Systolic Array Structured Pruning (SASP), a co-design strategy that combines systolic array acceleration and structured pruning with matching accelerator and tile size.
@@
  • We show how the insights collected from our framework enable the evaluation of figures of merit at different abstraction levels, including the assessment of QoS, performance, resource usage, and energy, as well as their trade-offs. We discuss how these can be effectively leveraged from the joint perspective of algorithmic optimization, system integration, and systolic array design.
  • Using a speech recognition case study, we show that SASP-based co-optimization of transformers and systolic arrays can lead to efficiency and speedup gains of up to 26% with minimal QoS impact.
II State of the Art
@@
By providing spatially-distributed computation with low control logic overhead, systolic arrays can effectively parallelize the execution of matrix multiplications, the dominant computing pattern in transformer inference [10]. To evaluate the benefits that systolic arrays can induce, SMAUG [15] and TiC-SAT [20] present system simulation infrastructures able to support complete inferences, including both their hardware-accelerated and their software-executed parts. These works showcase that even small-sized systolic arrays have the potential to reduce run-time by orders of magnitude. Performance is further improved when data is properly laid out in a tiled arrangement in memory according to the accelerator characteristics, in order to maximize spatio-temporal locality. Although such an approach has been adopted [11, 21], no attempt was made therein to prune computations as we do here with Systolic Array Structured Sparsity.
@@
Indeed, the ductility of DNN models makes them highly amenable to pruning. In the context of systolic array acceleration, pruning optimizations are categorized as either fine-grained or structured [16]. In the first case, specialized systolic arrays have been proposed which can leverage the presence of zero values in tiles by either clock gating processing elements [22] or by reordering operands [23, 24]. Nonetheless, fine-grained pruning requires a fair amount of control logic overhead in the accelerator design and impacts the regularity of data layout in memory, which may negate the intended benefits [25, 12].
@@
In this light, structured pruning strategies offer a promising alternative, as tiles of low-significance can be entirely skipped before processing them onto accelerators, when the tile size matches the target accelerator parallelism. Hence, speedups can be harnessed without requiring specialized hardware for sparsity management. This position has been adopted by previous works [26, 27, 28, 29, 30]. However, they only provide a partial view of the ensuing design space. In particular, [26, 28] and [27] adopt a system-level stance, exploring the potential for acceleration of co-designed pruning strategies and data-parallel accelerators, but do not investigate the impact on Quality of Service (QoS, e.g. accuracy, word error rate) of the performed pruning. Conversely, [29] and [30] provide an algorithmic-level assessment of the effect of structured sparsity, but neglect the hardware and architectural implication of adopting a matched accelerator design. To the best of our knowledge, a holistic view of the algorithmic-to-hardware space exposed by SASP is hence missing. Our paper aims at filling this gap.
III Co-designing Accelerators and Sparsity
@@
Figure 1: Qualitative radar plot illustrating two SASP solutions with different trade-offs: a slow and accurate one (red) employing a small accelerator and a low pruning rate, and a fast but inaccurate one (blue) using a large accelerator and a high pruning rate. Across all axes, higher is better.
Figure 2: Overview of Hardware-Software co-design framework.
@@
The dimensions of the design space exposed by Systolic Array Structured Pruning (SASP) solutions are illustrated in Figure 1. The illustrated figures of merit reside at widely different layers of the hardware/software stack. Hence, to enable co-optimization, we developed the integrated toolflow illustrated in Figure 2. The input to our framework is a trained transformer model and a target dataset, which, in our implementation, are defined via the ESPnet toolkit for automatic speech processing [3]. Hyper-parameters determine the size of the SASP tile (which sets the granularity of structured sparsity as well as the size of the systolic array) and the target sparsity rate. Moreover, both floating-point and weight-quantized implementations are supported.
@@
III-A Structured Pruning and Quantization
Figure 3: Tiled matrix multiplication with structured pruning.
We base our strategy on the observation that matrices employed by transformer models are much larger than the size of systolic arrays. Hence, operations to perform GEMM must be computed in a tiled fashion. As in [20], we herein consider a weight-stationary scenario, in which a tile of parameters is stored in the systolic array, and partial results are computed by streaming inputs/outputs to/from the accelerator. These are eventually aggregated via element-wise addition. As shown in Figure 3, in this setting, a tile containing only zero values can be completely skipped, saving both the time required to configure the systolic array and the time required to calculate the related partial results. In the example in Figure 3 it can be observed that the sparsity induced by the red weight tile lowers the workload required for the computation of the entire shaded column in the output.
We enforce structured sparsity by zeroing a percentage of tiles with the lowest L1-norm (sum of absolute values) across the entire model. This approach allows to heterogeneously prune GEMMs according to their sensitivity. In particular, feed-forward GEMMs are much more amenable to pruning than attention ones, so we focus on these for our exploration in Section IV-C.
After sparsification, post-training quantization can optionally be used to reduce the representation precision of weights from 32-bits floating point (FP32) to 8-bits integer (INT8). Finally, inference is performed on a target dataset, in order to gather QoS metrics such as Word Error Rate (WER).
III-B Full System Simulation
Run-time statistics on the deployment of the SASP-pruned models are collected in the gem5 [18] simulation environment, which allows specifying complex systems including hardware (processors, memory hierarchy) and software (operating system) components. To this end, we developed a systolic array gem5 module interfaced as a functional unit. Similarly to [20], the functional unit employs dedicated instructions, extending the ARM instruction set, to a) program weights, b) emulate the systolic array computation, and c) stream inputs/outputs (see Figure 4). We assume a 32-bit input-output interface, allowing to transfer one input and one output activation per custom instruction. As for weights, either a single FP32 or four INT8 values can be programmed in the array in a quantized or non-quantized setting, respectively.
@@ III-B Full System Simulation
III-C Systolic Array Architecture
@@
Figure 4: Architectural diagram of the systolic array, supporting FP32 activations and either non-quantized (FP32) or quantized (INT8) weights.
@@
Figure 5: Hardware diagram of the hybrid FP32_INT8 multiplier. This logic is bypassed in case any of the operands is equal to zero.
@@
IV Experimental Results
IV-A Setup
TABLE I: Parameters of the deployed ESPnet [3] model.
@@ IV-A Setup
layers
Feed-
forward
dimensions
Average
sequence
length
18 6 4 512 2 2048 128
TABLE II: Configuration of the simulated system.
@@
We evaluate our co-design approach on the LibriSpeech ASR corpus [19], using a transformer model implemented with ESPnet [3] using PyTorch [31], and targeting the encoder for SASP optimization, since its execution dominates run-time. The model structure corresponds to the parameters shown in Table I. The model is trained using 960 hours of the train set with speech perturbation (3$\times$ speed) [19] for 100 epochs and achieves 3.4% WER on both the development and test subsets (about 5 hours each). Further WER results in this paper are reported on the test subset. System simulations were run in the gem5-X variant [32] of the gem5 simulator [18]. We considered a single-core configuration having a 2-level cache hierarchy and running at 1 GHz, as detailed in Table II. Hardware syntheses of systolic array instances targeted the same 1 GHz timing constraint, and employed a TSMC 28nm technology node. Floating point arithmetic operators (adders and multipliers) were derived from the FPxx library using SpinalHDL [33], while the hybrid FP32_INT8 multiplier was implemented from scratch according to the design in Section III-C.
In all experiments, we considered systolic arrays of sizes ranging from $4\times 4$ to $32\times 32$. We spanned various structured pruning rates and investigated both FP32_FP32 and FP32_INT8 quantization schemes. Experiments at different abstraction tiers collected results on area, energy, run-time, and achieved Word Error Rate (WER). We detail each in the rest of the section. Then, we provide insights from a cross-tier point of view and summarize our findings in Section IV-F.
IV-B Hardware Exploration

### v2

@@ Abstract.
Efficient deployment of resource-intensive transformers on edge devices necessitates cross-stack optimization. We thus study the interrelation between structured pruning and systolic acceleration, matching the size of pruned blocks with the systolic array dimensions. In this setting, computations of pruned weight blocks can be skipped, reducing run-time and energy consumption, but potentially impacting quality of service (QoS). To evaluate the trade-offs between systolic array size and sparsity opportunities, we present a novel co-design framework that integrates algorithmic optimization, system simulation, and hardware design. Targeting speech recognition and machine translation using transformers as case study, we analyze how configuration choices across the stack affect performance metrics. Results demonstrate that structured pruning on systems featuring systolic array acceleration can effectively increase performance, while maintaining high QoS levels. Up to 44% system-wide speedups due to structured pruning and quantization were measured, with only 1.4% word error rate degradation on the standard LibriSpeech dataset.
@@
Systolic arrays, structured pruning, hardware-software co-design, edge AI.
journalyear: 2025copyright: rightsretainedconference: Great Lakes Symposium on VLSI 2025; June 30-July 2, 2025; New Orleans, LA, USAbooktitle: Great Lakes Symposium on VLSI 2025 (GLSVLSI ’25), June 30-July 2, 2025, New Orleans, LA, USAdoi: 10.1145/3716368.3735158isbn: 979-8-4007-1496-2/2025/06ccs: Hardware Hardware-software codesignccs: Computer systems organization Systolic arraysccs: Computing methodologies Neural networks
1. Introduction
@@
Transformers have fostered a revolution in machine learning, with applications ranging from classification (Dosovitskiy et al., 2020) to generative models for text and images (Brown et al., 2020), to speech recognition (Watanabe et al., 2018). However, their complex structure based on multiple attention and feed-forward layers (Vaswani et al., 2017) results in unprecedented computational requirements, posing significant challenges for their deployment. These are particularly acute in edge scenarios, where systems have to operate within constrained energy and performance envelopes.
@@
In this context, a plethora of optimization strategies have been proposed. On the software side (Chitty-Venkata et al., 2023), commonly used approaches involve reducing the precision of data representations (quantization) and removing parts that contribute the least to inference outcomes (pruning). As for hardware, efforts have mainly focused on the acceleration of the main computational kernel in transformers, i.e. General Matrix Multiplications (GEMMs). Although diverse solutions ranging from analog crossbars (Spoon et al., 2021)(Klein et al., 2022) to near-DRAM computing (Gómez-Luna et al., 2021)(Medina Morillas et al., 2024) have been investigated to this end, a particularly promising alternative is represented by systolic arrays (Xu et al., 2023). These two-dimensional meshes of processing elements can indeed parallelize the computation of a GEMM (or, more precisely, the computation of a GEMM tile), while presenting high parallelism, low resource requirements and only mandating a simple, low-overhead control logic.
@@
Recent works (Tuli et al., 2023; Yu et al., 2017; Jayakodi et al., 2021; Bringmann et al., 2021; Xi et al., 2020) have attempted to co-optimize software algorithms and hardware accelerators dedicated to transformer inference (Liang et al., 2021). Such a stance is particularly appealing at the crossroads of model pruning and systolic array acceleration. On the software side, pruning can be performed by eliding weights in regular block patterns (in a “structured” way) rather than as individual elements (Wen et al., 2016). While this approach introduces a constraint to pruning, and can hence result in lower overall sparsity rates, it substantially amplifies hardware-side optimization opportunities when matching the sizes of the pruned tile and the accelerator mesh. The exploration of this strategy, which we term Systolic Array Structured Pruning (SASP), is the focus of this work.
@@
To quantitatively explore these interrelations, we present a novel framework and methodology which, for the first time, integrates methods for a) the structured pruning and quantization of transformer algorithms, b) the system-level level modeling of accelerated systems executing them, and c) the hardware synthesis of accelerators. Our environment for SASP exploration builds on state-of-the-art methods for the training of transformers (ESPnet (Watanabe et al., 2018)) and for system simulation (gem5 (Binkert et al., 2011)). By employing a novel systolic array architectural template, it supports both floating point and weight-quantized data representations, as supported by ESPnet.
As a test case, we employ our exploration approach to analyze speech recognition (ASR) and machine translation (MT) applications, based on transformers processing the LibriSpeech (Panayotov et al., 2015) and MuST-C (Di Gangi et al., 2019) datasets. We observed that SASP can achieve, for a $32\times 32$ systolic array and a ESPnet ASR model trained on LibriSpeech, up to 44% speedup and 42% energy savings over a non-pruned, non-quantized system when employing a 20% pruning rate, resulting in a marginal Word Error Rate (WER) degradation of 1.4%.
@@
  • We perform a systematic exploration of Systolic Array Structured Pruning (SASP), a co-design strategy that combines systolic array acceleration and structured pruning with matching accelerator and tile size under quality of service constraints.
@@
  • We introduce a cross-stack framework to enable the evaluation of SASP through figures of merit at different abstraction levels, including the assessment of QoS, performance, resource usage, and energy, as well as their trade-offs. We discuss how these insights can be effectively leveraged from the joint perspective of algorithmic optimization, system integration, and systolic array design.
  • Across different models and tasks, we show that SASP-based co-optimization of transformers and systolic arrays can offer significant speedup and energy efficiency gains. For example, a joint ASR and MT case study achieves reductions in runtime and energy consumption of up to 51% and 34% while limiting QoS degradation to 4 BLEU points.
2. State of the Art
@@
By providing spatially-distributed computation with low control logic overhead, systolic arrays can effectively parallelize the execution of matrix multiplications, the dominant computing pattern in transformer inference (Xu et al., 2023). To evaluate the benefits that systolic arrays can induce, SMAUG (Xi et al., 2020) and TiC-SAT (Amirshahi et al., 2023) present system simulation infrastructures able to support complete inferences, including both their hardware-accelerated and their software-executed parts. These works showcase that even small-sized systolic arrays have the potential to reduce run-time by orders of magnitude. Performance is further improved when data is properly laid out in a tiled arrangement in memory according to the accelerator characteristics, in order to maximize spatio-temporal locality. Although such an approach has been adopted (Tuli et al., 2023; Amirshahi et al., 2024), no attempt was made therein to prune computations as we do here with Systolic Array Structured Sparsity.
@@
Indeed, the ductility of DNN models makes them highly amenable to pruning. In the context of systolic array acceleration, pruning optimizations are categorized as either fine-grained or structured (Liang et al., 2021). In the first case, specialized systolic arrays have been proposed which can leverage the presence of zero values in tiles by either clock gating processing elements (Liu et al., 2020) or by reordering operands (Tang et al., 2024; Lu et al., 2021). Nonetheless, fine-grained pruning requires a fair amount of control logic overhead in the accelerator design and impacts the regularity of data layout in memory, which may negate the intended benefits (Ma et al., 2021; Yu et al., 2017).
@@
In this light, structured pruning strategies offer a promising alternative, as tiles of low significance can be entirely skipped before processing them onto accelerators, when the tile size matches the target accelerator parallelism. Hence, speedups can be harnessed without requiring specialized hardware for sparsity management. This position has been adopted by previous works (Sun et al., 2023; Kang, 2019; Asgari et al., 2019; Li et al., 2020; Ben Letaifa and Rouas, 2023). However, they only provide a partial view of the ensuing design space. In particular, (Sun et al., 2023; Asgari et al., 2019) and (Kang, 2019) adopt a system-level stance, exploring the potential for acceleration of co-designed pruning strategies and data-parallel accelerators, but do not investigate the impact on quality of service (QoS, e.g. accuracy, word error rate) of the performed pruning. Conversely, (Li et al., 2020) and (Ben Letaifa and Rouas, 2023) provide an algorithmic-level assessment of the effect of structured sparsity, but neglect the hardware and architectural implication of adopting a matched accelerator design. To the best of our knowledge, a holistic analysis of the algorithmic-to-hardware space exposed by SASP is hence missing. Our paper aims at filling this gap.
3. Co-designing Accelerators and Sparsity
@@
Figure 1. Qualitative radar plot illustrating two SASP solutions with different trade-offs: a slow and accurate one (red) employing a small accelerator and a low pruning rate, and a fast but inaccurate one (blue) using a large accelerator and a high pruning rate. Across all axes, higher is better.
Figure 2. Overview of Hardware-Software co-design framework.
@@
The dimensions of the design space exposed by Systolic Array Structured Pruning (SASP) solutions are illustrated in Figure 1. The illustrated figures of merit reside at widely different layers of the hardware/software stack. Hence, to enable co-optimization, we developed the integrated toolflow illustrated in Figure 2. The input to our framework is a trained transformer model and a target dataset, which, in our implementation, are defined via the ESPnet toolkit for automatic speech processing (Watanabe et al., 2018). Hyper-parameters determine the size of the SASP tile (which sets the granularity of structured sparsity as well as the size of the systolic array) and the target sparsity rate. Moreover, both floating-point and weight-quantized implementations are supported.
@@
3.1. Structured Pruning and Quantization
Figure 3. Tiled matrix multiplication with structured pruning.
Matrices employed by transformer models are commonly much larger than the size of systolic arrays. Hence, operations to perform GEMM must be tiled. As in (Amirshahi et al., 2023), we herein consider a weight-stationary scenario, in which a tile of parameters is stored in the systolic array, and partial results are computed by streaming inputs/outputs to/from the accelerator. These are eventually aggregated via element-wise addition.
As shown in Figure 3, in this setting a tile containing only zero values can be completely skipped, saving both the time required to configure the systolic array and the time required to calculate the related partial results. In the example in Figure 3 it can be observed that the sparsity induced by the red weight tile lowers the workload required for the computation of the entire shaded column in the output. We enforce structured sparsity by zeroing a percentage of tiles with the lowest L1-norm (sum of absolute values) across the entire model. This approach allows to heterogeneously prune GEMMs according to their sensitivity. In particular, feed-forward GEMMs are much more amenable to pruning than attention ones, so we focus on these for our exploration in Section 4.3.
After sparsification, post-training quantization can optionally be used to reduce the representation precision of weights from 32-bits floating point (FP32) to 8-bits integer (INT8). Finally, inference is performed on a target dataset, in order to gather QoS metrics such as Word Error Rate (WER) or BLEU.
3.2. Full System Simulation
Run-time statistics on the deployment of the SASP-pruned models are collected in the gem5 (Binkert et al., 2011) simulation environment, which allows specifying complex systems including hardware (processors, memory hierarchy) and software (operating system) components. To this end, we developed a gem5 custom functional unit modeling a tighly coupled systolic array. Similarly to (Amirshahi et al., 2023), the functional unit employs dedicated instructions, extending the ARM instruction set, to a) program weights, b) perform the systolic array computation, and c) stream inputs/outputs (see Figure 4). We assume a 32-bit input-output interface, allowing to transfer one input and one output activation per custom instruction. As for weights, either a single FP32 or four INT8 values can be programmed in the array in a quantized or non-quantized setting, respectively.
@@ 3.2. Full System Simulation
3.3. Systolic Array Architecture
Figure 4. Architectural diagram of the systolic array, supporting FP32 activations and either non-quantized (FP32) or quantized (INT8) weights.
Figure 5. Hardware diagram of the hybrid FP32_INT8 multiplier. This logic is bypassed in case any of the operands is equal to zero.
@@ 3.3. Systolic Array Architecture
4. Experimental Results
4.1. Setup
Table 1. Parameters and Quality of Service (QoS) of the deployed models, targeting automatic speech recognition (ASR) and machine translation (MT). Training sets are augmented using 3$\times$ speed perturbation. QoS of the model on the MuST-C dataset (Di Gangi et al., 2019) are obtained for their execution in cascade (ASR $\rightarrow$ MT).
Toolkit Task Dataset
@@ 4.1. Setup
size
Training
set
Training
epochs
Test
set
Model
QoS
SASP
QoS
target
ESPnet ASR LibriSpeech 18 6 4 512 2048 960 h 100 5 h 3.5% WER 5% WER
ESPnet2 ASR LibriSpeech 12 6 8 512 2048 960 h 100 5 h 3.2% WER 5% WER
ESPnet2 ASR + MT MuST-C 18 / 6 6 / 6 4 / 4 128 / 128 2048 / 1024 560 h 35 / 200 5 h 31 BLEU 27 BLEU
Table 2. Configuration of the simulated system.
@@
To evaluate our framework’s capabilities, we focus on automatic speech recognition and machine translation, two challenging and commercially relevant edge AI tasks, using the LibriSpeech (Panayotov et al., 2015) and MuST-C (Di Gangi et al., 2019) corpora. Using ESPnet (Watanabe et al., 2018) and PyTorch (Paszke et al., 2019), we implemented transformer models targeting the encoder for SASP optimization, since its execution dominates run-time. The structure and training method of the models correspond to the parameters shown in Table 1. All the quality of service (QoS) results in this paper are reported on the respective corpora test sets and are comparable to state-of-the-art results (Guo et al., 2021). System simulations were run in the gem5-X variant (Qureshi et al., 2019) of the gem5 simulator (Binkert et al., 2011). We considered a single-core configuration having a 2-level cache hierarchy and running at 1 GHz, as detailed in Table 2. Hardware syntheses of systolic array instances targeted the same 1 GHz timing constraint, and employed a TSMC 28nm technology node. Floating point arithmetic operators (adders and multipliers) were derived from the FPxx library using SpinalHDL (Verbeure, [n. d.]), while the hybrid FP32_INT8 multiplier was implemented from scratch according to the design in Section 3.3.
In all experiments, we considered systolic arrays of sizes ranging from $4\times 4$ to $32\times 32$. We spanned various structured pruning rates and investigated both FP32_FP32 and FP32_INT8 quantization schemes. Experiments at different abstraction tiers collected results on area, energy, run-time, and achieved quality of service (QoS). We detail each in the rest of the section. Then, we provide insights from a cross-tier point of view and summarize our findings in Section 4.6.
4.2. Hardware Exploration